OpenStudy (anonymous):

Scores on a math midterm are distrubted under a standard normal curve. A randomly selected student had a score of 55, which represented a z score of -2.35. If the mean for the exam was 76, determine the standard deviation for the exam.

OpenStudy (kropot72):

\[z=\frac{X-\mu}{\sigma}\] \[-2.35=\frac{55-76}{\sigma}\] \[\sigma=\frac{55-76}{-2.35}\] Can you calculate sigma (the standard deviation) now?

OpenStudy (anonymous):

how do i do that?